# Spontaneous magnon decay in two-dimensional altermagnets

### v1

@@ Abstract
In the above example the convexity of the quasiparticle dispersion guarantees that the relevant decay process satisfies energy and momentum conservation. Conversely, quasiparticles with a concave dispersion cannot spontaneously decay to zero temperature, even if the ground state exhibits quantum fluctuations. This is the reason why magnons in quantum Heisenberg antiferromagnets with nearest-neighbor exchange on a square lattice are stable at zero temperature [4, 5]. In this work we show that in some sectors of the Brillouin zone such a kinematic constraint does not apply to magnons in the recently discovered altermagnets [6, 7, 8], so that magnons in altermagnets can in fact spontaneously decay. For simplicity we focus here on two-dimensional altermagnets and show that the corresponding decay rate is dominated by the decay of one magnon into three. For small momenta the decay rate is proportional to $k^{5}$ and depends on the direction in momentum space. Interestingly, magnon decay in altermagnets is found to be chirality selective in the sense that for fixed $\bm{k}$ only the magnons with one specific chirality can spontaneously decay while the magnons with the opposite chirality are stable. Spontaneous magnon decay in altermagnets has recently been discussed in Ref. [9]; however, an analytic understanding of the underlying physical mechanism has not been obtained.
@@
where $\bm{S}_{\bm{R}}$ are spin-$S$ operators localized at sites $\bm{R}$ of a square lattice with $N$ sites, lattice spacing $a$ and basis vectors $\bm{a}_{1}=a\hat{\bm{x}}$ and $\bm{a}_{2}=a\hat{\bm{y}}$. We have divided the lattice into two sublattices $A$ and $B$ as shown in Fig. 1. The antiferromagnetic exchange coupling $J>0$ connects all pairs of nearest-neighbor spins while the diagonal coupling $D$ connects spins across the diagonals $\bm{d}_{1}=a(\hat{\bm{x}}+\hat{\bm{y}})$ and $\bm{d}_{2}=a(-\hat{\bm{x}}+\hat{\bm{y}})$ on half of the plaquettes. On the diagonals of the other plaquettes the spins are connected by a different coupling $E$ such that the distinct plaquettes form a checkerboard pattern as illustrated in Fig. 1.
@@
where the sum is over the reduced Brillouin zone associated with the A-sublattice containing $N/2$ sites, $\alpha_{\bm{k}}$ and $\beta_{\bm{k}}$ are canonical magnon annihilation operators, and the dispersion of magnons with chirality $p=\pm$ is
@@
Figure 2: Magnon dispersions $\omega_{\bm{k}}^{\pm}$ defined in Eq. (3) for the path in the Brillouin zone shown in the inset where the M-point represents the wavevector $(0,\pi/a)$. The graph is for $D/J=-1/2$ and $E/J=-1/4$.
@@
Figure 3: (a): Graphical representation of the bosonized two-body Hamiltonian ${\cal{H}}_{4}$ of an altermagnet in terms of magnon operators. Here bold numbers $\bm{1},\bm{2},\bm{3},\bm{4}$ represent the lattice momenta $\bm{k}_{1}$, $\bm{k}_{2}$, $\bm{k}_{3}$, $\bm{k}_{4}$ which are conserved up a a vector $\bm{G}$ of the reciprocal lattice associated with the A-sublattice, i.e., $e^{i\bm{G}\cdot\bm{R}}=1$ for $\bm{R}\in A$. Outgoing/incoming blue arrows represent $\alpha^{\dagger}_{-\bm{k}}$/$\alpha_{\bm{k}}$, while outgoing/incoming red arrows represent $\beta^{\dagger}_{-\bm{k}}$/$\beta_{\bm{k}}$. Arrows entering or leaving a vertex at the same point indicate that the vertex is symmetric with respect to the exchange of the corresponding momenta. (b): Diagrammatic representation of the two-loop corrections to the magnon self-energies at zero temperature. Internal lines with arrows represent imaginary-frequency magnon propagators.
@@
Given the magnon Hamiltonian ${\cal{H}}_{2}+{\cal{H}}_{4}$ we may now calculate the magnon self-energies using the imaginary time formulation of diagrammatic perturbation theory. To obtain the leading contribution to the magnon damping, we have to calculate all diagrams up to second order in the interaction ${\cal{H}}_{4}$. Let us focus on the diagonal self-energy $\Sigma_{\alpha}(\bm{k},i\omega)$ of the $\alpha$-magnon, where $i\omega$ is a bosonic Matsubara frequency. Due to the rather complicated interaction ${\cal{H}}_{4}$ eight distinct second-order diagrams contribute to $\Sigma_{\alpha}(\bm{k},i\omega)$ which depend on various combinations of the interaction vertices. Fortunately, in the limit of vanishing temperature $T\rightarrow 0$ only the first diagram shown in Fig. 3 (b) survives. This diagram has the unique property that the arrows attached to all external and internal lines point into the same direction. After evaluating the relevant frequency sums at finite $T$ we find that the contribution of this diagram to the self-energy of the $\alpha$-magnon can be written as
@@
×$\times W(\bm{p},\bm{q};\bm{k})\frac{(1+n_{\bm{p}}^{+})(1+n_{\bm{q}}^{+})(1+n^{-}_{\bm{k}-\bm{p}-\bm{q}})}{i\omega-\omega^{+}_{\bm{p}}-\omega^{+}_{\bm{q}}-\omega^{-}_{\bm{k}-\bm{p}-\bm{q}}},$ (11)
@@
Figure 4: (a) For small non-linearity energy and momentum conservation imply that the decay of one magnon with momentum $\bm{k}$ into three magnons with momenta $\bm{p}$, $\bm{q}$, and $\bm{k}-\bm{p}-\bm{q}$ is only possible if $\bm{p}$ and $\bm{q}$ are almost aligned with $\bm{k}$. (b) Elliptic scattering surface in the $p_{\bot}$-$q_{\bot}$ plane associated with spontaneous magnon decay. The red dashed line represents the solution of Eq. (15) which is valid for small $ka$ and to leading order in $p_{\bot}/k$ and $q_{\bot}/k$. The yellow curve is obtained from the numerical solution of the original Eq. (14). We have chosen the following parameters: $D/J=-1/2$, $E/J=-1/4$, $k_{x}a=k_{y}a=0.003\pi/\sqrt{2}$, $p_{\parallel}a=q_{\parallel}a=0.001\pi$. (c) For larger wavevectors the shape of the scattering surface deviates from an ellipse.
@@
Keeping in mind that momentum conservation requires $0<p_{\parallel}<k$, $0<q_{\parallel}<k$, and $0<k-p_{\parallel}-q_{\parallel}<k$, the quadratic form on the left-hand side of Eq. (15) is positive definite. For $m>0$ this implies that Eq. (15) has only a solution if $\hat{k}_{x}\hat{k}_{y}>0$, i.e. $\bm{k}$ must lie in the first or in the third quadrant of the Brillouin zone. For fixed $p_{\parallel}$ and $q_{\parallel}$ the solution of Eq. (15) is then an ellipse in the $p_{\bot}$-$q_{\bot}$ plane, as illustrated in Fig. 4 (b). The elliptic approximation to the scattering surface is accurate to leading order in $ka$ and $k/|mc|$. For larger wavevectors the shape of the scattering surfaces deviates from an ellipse, as illustrated in Fig. 4 (c).
@@
Note that in this limit the vertex is independent of the altermagnetic couplings $D$ and $E$. Keeping in mind that for the scattering geometry of interest the loop momenta $\bm{p}$ and $\bm{q}$ are almost parallel to the external momentum $\bm{k}$, the relevant squared matrix element can be approximated by
@@
Substituting this into Eq. (13) and using the elliptic approximation (15) for the scattering surface, we may scale out the momentum dependence by setting $p_{\parallel}=kx_{1}$, $q_{\parallel}=kx_{2}$, and $p_{\bot}=\sqrt{2k^{3}/|mc|}y_{1}$, $q_{\bot}=\sqrt{2k^{3}/|mc|}y_{2}$. Note that the transverse momenta scale as $k^{3/2}$, whereas the longitudinal momenta are of order $k$ so that for sufficiently small $k$ our assumption that $p_{\bot}$ and $q_{\bot}$ are small compared with $p_{\parallel}$ and $q_{\parallel}$ is justified. The transverse integration over $y_{1}$ and $y_{2}$ can then be carried out analytically, while the remaining two-dimensional integral over $x_{1}$ and $x_{2}$ can be evaluated numerically. Our final result for the decay rate of $\alpha$-magnons is
@@
where $\Theta(x)=1$ for $x>0$ and $\Theta(x)=0$ for $x<0$. The decay rate of the $\beta$-magnons can be obtained by substituting $\Theta(k_{x}k_{y}/m)\rightarrow\Theta(-k_{x}k_{y}/m)$ in Eq. (18). We conclude that spontaneous magnon decay in two-dimensional altermagnets is chirality selective: for positive $m$ the $\alpha$-magnons can spontaneously decay only in the first and third quadrant of the Brillouin zone, while the $\beta$-magnons can decay only in the second and fourth quadrant. Moreover, the decay rates are maximal for momenta parallel to the diagonals of the Brillouin zone and vanish along the symmetry axes of the crystal, as illustrated in Fig. 5.
Figure 5: Illustration of the chirality selectivity of the spontaneous magnon decay rate in two-dimensional altermagnets for small momenta. The intensity of the blue shading is proportional to the spontaneous decay rate $\gamma_{\alpha}(\bm{k})$ of the $\alpha$-magnon given in Eq. (18), while the red shading represents the corresponding decay rate $\gamma_{\beta}(\bm{k})$ of the $\beta$-magnon. We only show the long-wavelength regime $|\bm{k}|\leq 1/a$ where the approximations leading to Eq. (18) are expected to be accurate.
Our result (18) should be compared with the Beliaev damping of sound in the ground state of the superfluid Bose gas, which is proportional $k^{3}$ in two dimensions [3]. In Eq. (18) two extra powers of $k$ are generated by the strong momentum dependence of the squared scattering vertex in Eq. (17).
@@
Note added: While preparing this manuscript we have become aware of independent complementary work on magnon decays in general models for two-dimensional altermagnets by Eto et al. [18], who also presented numerical results for arbitrary momenta.

### v2

@@ I Introduction
In the above example the convexity of the quasiparticle dispersion guarantees that the relevant decay process satisfies energy and momentum conservation. Conversely, quasiparticles with a concave dispersion cannot spontaneously decay to zero temperature, even if the ground state exhibits quantum fluctuations. This is the reason why magnons in quantum Heisenberg antiferromagnets with nearest-neighbor exchange on a square lattice are stable at zero temperature [4, 5]. In this work we show that in some sectors of the Brillouin zone such a kinematic constraint does not apply to magnons in the recently discovered altermagnets [6, 7, 8], so that magnons in altermagnets can in fact spontaneously decay. For simplicity we focus here on two-dimensional altermagnets and show that the corresponding decay rate is dominated by the decay of one magnon into three. For small momenta the decay rate is proportional to $k^{5}$ and depends on the direction in momentum space [9]. Interestingly, magnon decay in altermagnets is found to be chirality selective in the sense that for fixed $\bm{k}$ only the magnons with one specific chirality can spontaneously decay while the magnons with the opposite chirality are stable. Spontaneous magnon decay in altermagnets has recently been discussed in Ref. [11]; however, an analytic understanding of the underlying physical mechanism has not been obtained.
II Minimal spin model
@@
[…]
@@
where the sum is over the reduced Brillouin zone associated with the $A$-sublattice containing $N/2$ sites, $\alpha_{\bm{k}}$ and $\beta_{\bm{k}}$ are canonical annihilation operators for magnons with chirality $p=\pm$ and dispersions
@@
Figure 2: Magnon dispersions $\omega_{\bm{k}}^{\pm}$ defined in Eq. (3) for the path in the Brillouin zone shown in the inset where the M-point represents the wavevector $(0,\pi/a)$. The graph is for D/J=1/2D/J=-1/2 and E/J=1/4E/J=-1/4.
@@ III Spontaneous magnon decay
Figure 3: (a): Graphical representation of the bosonized two-body Hamiltonian ${\cal{H}}_{4}$ of an altermagnet in terms of magnon operators. Here bold numbers $\bm{1},\bm{2},\bm{3},\bm{4}$ represent the lattice momenta $\bm{k}_{1}$, $\bm{k}_{2}$, $\bm{k}_{3}$, $\bm{k}_{4}$ which are conserved up to a vector $\bm{G}$ of the reciprocal lattice associated with the AA-sublattice, i.e., $e^{i\bm{G}\cdot\bm{R}}=1$ for $\bm{R}\in A$. Outgoing/incoming blue arrows represent $\alpha^{\dagger}_{-\bm{k}}$$\alpha_{\bm{k}}$, while outgoing/incoming red arrows represent $\beta^{\dagger}_{-\bm{k}}$$\beta_{\bm{k}}$. Arrows entering or leaving a vertex at the same point indicate that the vertex is symmetric with respect to the exchange of the corresponding momenta. (b): Diagrammatic representation of the two-loop corrections to the magnon self-energies at zero temperature. Internal lines with arrows represent imaginary-frequency magnon propagators.
@@
Given the magnon Hamiltonian ${\cal{H}}_{2}+{\cal{H}}_{4}$ we may now calculate the magnon self-energies using the imaginary time formulation of diagrammatic perturbation theory. To obtain the leading contribution to the magnon damping, we have to calculate all diagrams up to second order in the interaction ${\cal{H}}_{4}$. Let us focus on the diagonal self-energy $\Sigma_{\alpha}(\bm{k},i\omega)$ of the $\alpha$-magnon with chirality $p=+$, where $i\omega$ is a bosonic Matsubara frequency. Due to the rather complicated interaction ${\cal{H}}_{4}$ eight distinct second-order diagrams contribute to $\Sigma_{\alpha}(\bm{k},i\omega)$ which depend on various combinations of the interaction vertices. Fortunately, in the limit of vanishing temperature $T\rightarrow 0$ only the first diagram shown in Fig. 3 (b) survives. This diagram has the unique property that the arrows attached to all external and internal lines point into the same direction. After evaluating the relevant frequency sums at finite $T$ we find that the contribution of this diagram to the self-energy of the $\alpha$-magnon can be written as
@@
$\times W(\bm{p},\bm{q};\bm{k})\frac{(1+n_{\bm{p}}^{+})(1+n_{\bm{q}}^{+})(1+n^{-}_{\bm{k}-\bm{p}-\bm{q}})}{i\omega-\omega^{+}_{\bm{p}}-\omega^{+}_{\bm{q}}-\omega^{-}_{\bm{k}-\bm{p}-\bm{q}}},$ (11)
@@
Figure 4: (a) For small non-linearity energy and momentum conservation imply that the decay of one magnon with momentum $\bm{k}$ into three magnons with momenta $\bm{p}$, $\bm{q}$, and $\bm{k}-\bm{p}-\bm{q}$ is only possible if $\bm{p}$ and $\bm{q}$ are almost aligned with $\bm{k}$. (b) Elliptic scattering surface in the $p_{\bot}$-$q_{\bot}$ plane associated with spontaneous magnon decay. The red dashed line represents the solution of Eq. (15) which is valid for small kaka and to leading order in $p_{\bot}/k$ and $q_{\bot}/k$. The yellow curve is obtained from the numerical solution of the original Eq. (14). We have chosen the following parameters: D/J=1/2D/J=-1/2, E/J=1/4E/J=-1/4, $k_{x}a=k_{y}a=0.003\pi/\sqrt{2}$, $p_{\parallel}a=q_{\parallel}a=0.001\pi$. (c) For larger wavevectors the shape of the scattering surface deviates from an ellipse.
@@
Keeping in mind that momentum conservation requires $0<p_{\parallel}<k$, $0<q_{\parallel}<k$, and $0<k-p_{\parallel}-q_{\parallel}<k$, the quadratic form on the left-hand side of Eq. (15) is positive definite. For $m>0$ this implies that Eq. (15) has only a solution if $\hat{k}_{x}\hat{k}_{y}>0$; i.e., $\bm{k}$ must lie in the first or in the third quadrant of the Brillouin zone. For fixed $p_{\parallel}$ and $q_{\parallel}$ the solution of Eq. (15) is then an ellipse in the $p_{\bot}$-$q_{\bot}$ plane, as illustrated in Fig. 4 (b). The elliptic approximation to the scattering surface is accurate to leading order in $ka$ and $k/|mc|$. For larger wavevectors the shape of the scattering surfaces deviates from an ellipse, as illustrated in Fig. 4 (c).
@@
Note that in this limit the vertex is independent of the altermagnetic couplings $D$ and $E$, so that magnon decay in altermagnets is due to the specific kinematics of the band splitting. Keeping in mind that for the scattering geometry of interest the loop momenta $\bm{p}$ and $\bm{q}$ are almost parallel to the external momentum $\bm{k}$, the relevant squared matrix element can be approximated by
@@
Substituting this into Eq. (13) and using the elliptic approximation (15) for the scattering surface, we may scale out the momentum dependence by setting $p_{\parallel}=kx_{1}$, $q_{\parallel}=kx_{2}$, and $p_{\bot}=\sqrt{2k^{3}/|mc|}y_{1}$, $q_{\bot}=\sqrt{2k^{3}/|mc|}y_{2}$. Note that the transverse momenta scale as $k^{3/2}$, whereas the longitudinal momenta are of order $k$ so that for sufficiently small $k$ our assumption that $p_{\bot}$ and $q_{\bot}$ are small compared with $p_{\parallel}$ and $q_{\parallel}$ is justified. The transverse integration over $y_{1}$ and $y_{2}$ can then be carried out analytically, while the remaining two-dimensional integral over $x_{1}$ and $x_{2}$ can be evaluated numerically. Our final result for the decay rate of $\alpha$-magnons is [17]
@@
where $\Theta(x)=1$ for $x>0$ and $\Theta(x)=0$ for $x<0$. The decay rate of the $\beta$-magnons with chirality $p=$- can be obtained by substituting $\Theta(k_{x}k_{y}/m)\rightarrow\Theta(-k_{x}k_{y}/m)$ in Eq. (18).
IV Conclusions
We conclude that spontaneous magnon decay in two-dimensional altermagnets is chirality selective: for positive $m$ the $\alpha$-magnons can spontaneously decay only in the first and third quadrant of the Brillouin zone, while the $\beta$-magnons can decay only in the second and fourth quadrant. Moreover, the decay rates are maximal for momenta parallel to the diagonals of the Brillouin zone and vanish along the symmetry axes of the crystal, as illustrated in Fig. 5.
Figure 5: Illustration of the chirality selectivity of the spontaneous magnon decay rate in two-dimensional altermagnets for small momenta. The intensity of the blue shading is proportional to the spontaneous decay rate $\gamma_{\alpha}(\bm{k})$ of the $\alpha$-magnon given in Eq. (18), while the red shading represents the corresponding decay rate $\gamma_{\beta}(\bm{k})$ of the $\beta$-magnon. We only show the long-wavelength regime $|\bm{k}|\ll 1/a$ where the approximations leading to Eq. (18) are expected to be accurate.
Our result (18) should be compared with the Beliaev damping of sound in the ground state of the superfluid Bose gas, which is proportional $k^{3}$ in two dimensions [3]. In Eq. (18) two extra powers of $k$ are generated by the strong momentum dependence of the squared scattering vertex in Eq. (17). This is similar to the XXZ ferromagnet [18], where the hydrodynamic momentum dependence of the scattering vertex generates four extra powers of $k$, so that the damping behaves as $k^{7}$ at long wavelengths. Note also that our result (18) for the damping at long wavelengths is not specific to the checkerboard model (1), but generic to all models with the same symmetries, since we expressed it in terms of the long-wavelength parameters $c$ and $m$.